# Hermitian-symplectic and Kähler structures on degenerate twistor deformations

### v1

@@ 1.1 Degenerate twistor deformations and the Tate–Shafarevich twist
In [Mar], Markman has studied the Tate–Shafarevich twist of a Lagrangian fibration on a Hilbert scheme of a K3 surface. Markman calls a compact holomorphically symplectic manifold $(M,\Omega)$ “M-special” if the space $\langle[\operatorname{Re}\Omega],[\operatorname{Im}\Omega]\rangle\subset H^{2}(M,{\mathbb{R}})$ does not contain non-zero rational cohomology classes. He has shown that the Tate–Shafarevich twist of a Lagrangian fibration on a Hilbert scheme of a K3 surface is Kähler, if it is not M-special. In [AR], this result was generalized to all Lagrangian fibrations without multiple fibers in codimension 1. In the present paper we give an independent proof that works for arbitrary compact hyperkähler manifolds without the assumption of non-M-speciality.
@@ 1.2 Non-Kähler deformations of hyperkähler manifolds
The earliest publication on Kählerness of holomorphically symplectic manifolds was a preprint of Andrey Todorov [T]. Using a Harvey-Lawson-type Hahn-Banach argument, Todorov tried to show that a simply connected, compact, holomorphically symplectic manifold with $H^{2,0}(M)={\mathbb{C}}$ is Kähler.
@@
In 4.2, we prove this conjecture when the central fiber $M_{0}$ is Hermitian symplectic. In [Per], A. Perego proved 1.2 when $M_{0}$ satisfies $b_{2}(M_{0})=h^{2,0}(M_{0})+h^{1,1}(M_{0})+h^{0,2}(M_{0})$.
@@
In [Po], this conjecture is proven when $M_{t}$, $t\neq 0$ are projective, and $h^{0,1}(M_{0})=h^{0,1}(M_{t})$ for $t\neq 0$; in this case, the central fiber is Moishezon.
@@ 1.3 Hermitian symplectic geometry and Streets–Tian conjecture
Theorem 1.4: Let $\operatorname{\sf Teich}$ be the Teichmuller space of holomorphically symplectic complex structures, and $I\in\operatorname{\sf Teich}$ a holomorphically symplectic complex structure obtained as a limit $I=\lim_{i}I_{i}$, where $I_{i}\in\operatorname{\sf Teich}$ are holomorphically symplectic complex structures of Kähler type. Assume that $(M,I)$ is Hermitian symplectic. Then $(M,I)$ is also of Kähler type.
@@ 2.1 Hyperkähler manifolds
Consider the collection $\operatorname{\sf Comp}(M)$ of complex structures of hyperkähler type on $M$ and the group $\operatorname{\sf Diff}^{\circ}(X)$ of diffeomorphisms of $M$ isotopic to the identity. The quotient $\operatorname{\sf Teich}(M)=\operatorname{\sf Comp}(M)/\operatorname{\sf Diff}^{\circ}(M)$ carries the structure of a non-Hausdorff complex manifold and is called the Teichmüller space of $M$. The globall Torelli theorem for hyperkähler manifolds [V1] describes the Teichmüller space in terms of the cohomology of $M$. Define the period domain:
@@ 2.2 Degenerate twistor deformations
The periods of the manifolds $\mathcal{M}_{t}$ form an affine line in the period domain. More precisely, consider the subspace $W\subset H^{2}(M,\mathbb{C})$ spanned by $\Omega$, $\overline{\Omega}$ and $\pi^{*}\eta$. Then the restriction of the BBF form $q|_{W}$ is a degerate positive-semidefinite quadratic form with one-dimensional kernel spanned by $\pi^{*}\eta$. The corresponding quadric $Q_{W}$ is the union of two projective lines intersecting at the point $[\pi^{*}\eta]$. The complement to that point in $Q_{W}$ is the disjoint union of two affine lines, one of them passing through $[\Omega]$ and the other through $[\overline{\Omega}]$. The first of the two affine lines is formed exactly by the periods of the manifolds $\mathcal{M}_{t}$, $t\in\mathbb{C}$ of the degenerate twistor family. The other affine line corresponds to the complex-conjugate degenerate twistor family which one obtains replacing $\Omega$ by $\overline{\Omega}$.
@@ 2.3 Weakly positive and strongly positive forms
Definition 2.5: Let $(V,I)$ be a $2n$-dimensional real vector space eqipped with a complex structure operator $I$, $I^{2}=-\operatorname{Id}$. A weakly positive $(p,p)$-form on $V$ is a real $(p,p)$-form $\eta\in\Lambda^{p,p}_{\mathbb{R}}(V^{*})$ which satisfies
@@ 2.4 Positive currents
3 Hermitian symplectic structures and degenerate twistor deformations
@@
Let $V=\Lambda^{2}(M,{\mathbb{R}})$ be the Fréchet space of smooth two-forms equipped with the $C^{\infty}$-topology, $A$ the cone of two-forms with strictly positive $(1,1)$-part, and $W\subset V$ the space of closed two-forms. The intersection $A\cap W$ is the set of all Hermitian symplectic forms. If this set is empty, we use 3 to find a non-zero functional $\theta\in V^{*}$ which is positive on $A$ (this means that $\theta$ is a positive $(n-1,n-1)$-current) and vanising on $W$. The latter condition means that $\theta$ is exact. Indeed, $\langle\theta,du\rangle=0$ for all $u\in\Lambda^{1}M$ implies that $\theta$ is closed. It is exact, because otherwise it is non-zero on $W$ by Poincaré duality (the de Rham cohomology of currents is equal to the de Rham cohomology of differential forms, [De]).
@@ 4.1 Barlet spaces on commendable manifolds
4.2 Non-separated points in the Teichmüller space of Hermitian symplectic holomorphically symplectic manifolds
@@
The idea now is to take the graph $\Gamma_{t}$ of the bimeromorphism $\mathcal{X}_{t}\dashrightarrow\mathcal{M}_{t}$ and use 4.1 to show that there is a subsequence $\{\Gamma_{t_{i}}\}$ converging to a subvariety $\Gamma_{t_{0}}$ in $\mathcal{X}_{t_{0}}\times\mathcal{M}_{t_{0}}$. Then we use a topological argument to show that one of the irreducivle components of $\Gamma_{t_{0}}$ is a graph of a bimeromorphism $\mathcal{X}_{t_{0}}\dashrightarrow\mathcal{M}_{t_{0}}$.
@@
Denote by $p\colon X\times M\to X$ and $q\colon X\times M\to M$ the two projections. Let us consider the action of the correspondence given by $[Z]$ on the cohomology class of the symplectic form $\Omega$. Recall that by our assumptions $[Z]_{*}[\Omega]=[\Omega^{\prime}]$. For a connected component $Z_{i}$ of the cycle $Z$ denote by $\widetilde{Z}_{i}$ its resolution of singularities and by $\widetilde{p}_{i}\colon\widetilde{Z}_{i}\to X$ and $\widetilde{q}_{i}\colon\widetilde{Z}_{i}\to M$ the induced maps. Then $[Z_{i}]_{*}[\Omega]$ is the cohomology classs of the current $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$. For $i\neq i_{0}$ the map $\widetilde{p}_{i}$ is not dominant, and since $\widetilde{q}_{i}^{*}\Omega$ is a holomorphic two-form, the current $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$ vanishes. Indeed, a closed (2,0)-current is holomorphic, hence smooth by elliptic reguarity, but the support of $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$ is contained in a proper subvariety. Now consider the case $i=i_{0}$. Then $\widetilde{p}_{i_{0}*}(\widetilde{q}_{i_{0}}^{*}\Omega)$ is a holomorphic two-form on $X$, therefore it must be proportional to $\Omega^{\prime}$, because $X$ is hyperkähler of maximal holonomy. It follows that $\widetilde{q}_{i_{0}}^{*}\Omega$ is generically a symplectic form, so $\widetilde{q}_{i_{0}}$ must be dominant, hence $i_{0}=i_{1}$.
@@ 4.3 Degenerate twistor deformations are Kähler
Proof:  Denote by $U\subset\mathbb{A}^{1}$ the open subset of points of the base such that for $t\in U$ the fiber $\mathcal{M}_{t}$ is Kähler. Assume that $U\neq\mathbb{A}^{1}$ and let $x\in\partial U$. Then by 4.2, the fiber $\mathcal{M}_{t}$ is of Fujiki class C. Moreover, by 3 the manifold $\mathcal{M}_{t}$ admits a Hermitian symplectic form $\omega$. Then $\omega^{1,1}$ is a Hermitian metric and $dd^{c}(\omega^{1,1})=0$, therefore $\mathcal{M}_{t}$ is Kähler by [Ch, Theorem 0.2]  [^2] We thank the participants of the mathoverflow discussion for this reference or 5. It follows that $U=\mathbb{A}^{1}$.
5 Appendix: Hermitian symplectic forms and Fujiki class C symplectic manifolds
For the sake of completeness, we give a self-contained proof of [Ch, Theorem 0.2] for holomophically symplectic manifolds.
Theorem 5.1:  Assume that a compact holomorphically symplectic manifold $M$ is bimeromorphic to a hyperkähler manifold of maximal holonomy and that $\omega\in\Lambda^{2}M$ is a Hermitian symplectic form. Then $M$ is Kähler.
Proof:  Denote by $\Omega$ the holomorphic symplectic form on $M$ and let $f\colon M\dashrightarrow M^{\prime}$ be a bimeromorphic map to a hyperkähler manifold of maximal holonomy $M^{\prime}$. Then $\Omega^{\prime}=f_{*}\Omega$ is a holomorphic symplectic form on $M^{\prime}$. Since $M$ is of Fujiki class C, we may apply Hodge theory to it. Since $f$ maps a holomorphic volume form on $M$ to a holomorphic volume form on $M^{\prime}$, the indeterminacy locus of $f$ has codimension at least two. Hence $f_{*}\colon H^{2}(M,\mathbb{C})\to H^{2}(M^{\prime},\mathbb{C})$ is an isomorphism of Hodge structures. We will denote by $q$ the BBF form on $H^{2}(M^{\prime},\mathbb{C})$.
Let us denote $a=[\omega]^{1,1}$, the $(1,1)$-part of the cohomology class of the Hermitian symplectic form on $M$. Assume that $\omega^{\prime}\in\Lambda^{1,1}M^{\prime}$ is a Kähler form and $b=[\omega^{\prime}]\in H^{2}(M^{\prime},\mathbb{C})$ is its cohomology class. Let us observe that $q(f_{*}a,b)>0$. Indeed, resolving the indeterminacy of $f$ we can construct a manifold $X$ and two bimeromorphic maps $p_{1}\colon X\to M$, $p_{2}\colon X\to M^{\prime}$ such that $f=p_{2}\circ p_{1}^{-1}$. Then using the Fujiki relations for $q$ we get:
$q(f_{*}a,b)=\frac{q(b)}{\int_{M^{\prime}}b^{2n}}\int_{M^{\prime}}(f_{*}a)b^{2n-1}=\frac{q(b)}{\int_{M^{\prime}}b^{2n}}\int_{X}p_{1}^{*}(\omega^{1,1})\wedge(p_{2}^{*}\omega^{\prime})^{2n-1}>0,$
the latter inequality being valid because both $\omega^{1,1}$ and $\omega^{\prime}$ are strictly postitive forms. We conclude that the class $f_{*}a$ lies in the positive cone of $M^{\prime}$. It is clear from the definition that $\omega$ will stay Hermitian symplectic under small perturbations. Therefore, after possibly modifying $\omega$, we may assume that $f_{*}a$ lies in one of the Kähler chambers of $M^{\prime}$. Replacing $M^{\prime}$ by another bimeromorphic hyperkähler model we may assume that $f_{*}a$ is Kähler. The claim now follows from [Hu2, Proposition 2.1] which shows that the bimeromorphic map $f$ taking a Kähler class to a Kähler class is regular, hence defines an isomorphism of $M$ and $M^{\prime}$.
Acknowledgements: We are grateful to Ekaterina Amerik, Alexey Golota, Olivier Martin and Eyal Markman for interesting discussions of the subject.

### v2

@@ 1.1 Degenerate twistor deformations and the Tate–Shafarevich twist
In [Mar], Markman has studied the Tate–Shafarevich twist of a Lagrangian fibration on a Hilbert scheme of a K3 surface. Markman calls a compact holomorphically symplectic manifold $(M,\Omega)$ “M-special” if the space $\langle[\operatorname{Re}\Omega],[\operatorname{Im}\Omega]\rangle\subset H^{2}(M,{\mathbb{R}})$ does not contain non-zero rational cohomology classes. He has shown that the Tate–Shafarevich twist of a Lagrangian fibration on a Hilbert scheme of a K3 surface is Kähler, if it is not M-special. In [AR], this result was generalized to all Lagrangian fibrations without multiple fibers in codimension 1. In the present paper we give an independent proof that works for arbitrary compact hyperkähler manifolds without the assumption of non-M-specialty.
@@ 1.2 Non-Kähler deformations of hyperkähler manifolds
The earliest publication on Kählerness of holomorphically symplectic manifolds was a preprint of Andrey Todorov [T2]. Using a Harvey-Lawson-type Hahn-Banach argument, Todorov tried to show that a simply connected, compact, holomorphically symplectic manifold with $H^{2,0}(M)={\mathbb{C}}$ is Kähler.
@@
In 4.2, we prove this conjecture when the central fiber $M_{0}$ is Hermitian symplectic. In [Per], A. Perego discussed 1.2 when $M_{0}$ satisfies $b_{2}(M_{0})=h^{2,0}(M_{0})+h^{1,1}(M_{0})+h^{0,2}(M_{0})$.
@@
The application of this argument often brings an error, which was observed by Y.-T. Siu [S1, S2] when he was discussing a paper of A. Todorov [T1] “Applications of the Kähler-Einstein-Calabi-Yau metric to moduli of K3 surfaces.” The same problem occurs later in two important and very interesting papers [Per, A].
As explained by Siu, the gaps occur in the application of E. Bishop’s theorem on limits of complex analytic subvarieties. Suppose that we have two families of holomorphically symplectic manifolds over a disk, isomorphic except in the central fiber; as in Huybrechts’s proof, one considers the limit of the graphs $\Gamma_{t}$ of the isomorphisms between the non-central fibers. The idea is to apply Bishop’s theorem to show that this limit is a complex analytic subvariety with the same fundamental class.
To apply Bishop’s theorem, bounding the volume of the limiting cycle, one needs to choose a Hermitian form on the central fiber, such that the volume of $\Gamma_{t}$ with respect to this form stays bounded for all $t$. Even if we construct a sequence of Kähler metrics on the neighboring fibers whose cohomology classes converge to a cohomology class on the central fiber (such as in [Per, Lemma 2.5] and in [A, Lemma 4.6.1]), this does not imply the volume estimates. Indeed, a cohomology class with positive BBF square on a holomorphic symplectic manifold that is a limit of Kähler classes on neighboring fibers needs not be Kähler itself.
@@
In [Po], this conjecture is proven when $M_{t}$, $t\neq 0$ are projective, and $h^{0,1}(M_{0})=h^{0,1}(M_{t})$ for $t\neq 0$; in this case, the central fiber is Moishezon. Later, this result was amplified by [LRWW], who proved that a limit of Moishezon manifolds is Moishezon.
@@ 1.3 Hermitian symplectic geometry and Streets–Tian conjecture
Theorem 1.4: Let $\operatorname{\sf Teich}$ be the Teichmüller space of holomorphically symplectic complex structures, and $I\in\operatorname{\sf Teich}$ a holomorphically symplectic complex structure obtained as a limit $I=\lim_{i}I_{i}$, where $I_{i}\in\operatorname{\sf Teich}$ are holomorphically symplectic complex structures of Kähler type. Assume that $(M,I)$ is Hermitian symplectic. Then $(M,I)$ is also of Kähler type.
@@ 2.1 Hyperkähler manifolds
Consider the collection $\operatorname{\sf Comp}(M)$ of complex structures of hyperkähler type on $M$ and the group $\operatorname{\sf Diff}^{\circ}(X)$ of diffeomorphisms of $M$ isotopic to the identity. The quotient $\operatorname{\sf Teich}(M)=\operatorname{\sf Comp}(M)/\operatorname{\sf Diff}^{\circ}(M)$ carries the structure of a non-Hausdorff complex manifold and is called the Teichmüller space of $M$. The global Torelli theorem for hyperkähler manifolds [V1] describes the Teichmüller space in terms of the cohomology of $M$. Define the period domain:
@@ 2.2 Degenerate twistor deformations
The periods of the manifolds $\mathcal{M}_{t}$ form an affine line in the period domain. More precisely, consider the subspace $W\subset H^{2}(M,\mathbb{C})$ spanned by $\Omega$, $\overline{\Omega}$ and $\pi^{*}\eta$. Then the restriction of the BBF form $q|_{W}$ is a degenerate positive-semidefinite quadratic form with one-dimensional kernel spanned by $\pi^{*}\eta$. The corresponding quadric $Q_{W}$ is the union of two projective lines intersecting at the point $[\pi^{*}\eta]$. The complement to that point in $Q_{W}$ is the disjoint union of two affine lines, one of them passing through $[\Omega]$ and the other through $[\overline{\Omega}]$. The first of the two affine lines is formed exactly by the periods of the manifolds $\mathcal{M}_{t}$, $t\in\mathbb{C}$ of the degenerate twistor family. The other affine line corresponds to the complex-conjugate degenerate twistor family which one obtains replacing $\Omega$ by $\overline{\Omega}$.
@@ 2.3 Weakly positive and strongly positive forms
Definition 2.5: Let $(V,I)$ be a $2n$-dimensional real vector space equipped with a complex structure operator $I$, $I^{2}=-\operatorname{Id}$. A weakly positive $(p,p)$-form on $V$ is a real $(p,p)$-form $\eta\in\Lambda^{p,p}_{\mathbb{R}}(V^{*})$ which satisfies
@@ 2.4 Positive currents
3 Hermitian symplectic structures
@@
Let $V=\Lambda^{2}(M,{\mathbb{R}})$ be the Fréchet space of smooth two-forms equipped with the $C^{\infty}$-topology, $A$ the cone of two-forms with strictly positive $(1,1)$-part, and $W\subset V$ the space of closed two-forms. The intersection $A\cap W$ is the set of all Hermitian symplectic forms. If this set is empty, we use 3 to find a non-zero functional $\theta\in V^{*}$ which is positive on $A$ (this means that $\theta$ is a positive $(n-1,n-1)$-current) and vanishing on $W$. The latter condition means that $\theta$ is exact. Indeed, $\langle\theta,du\rangle=0$ for all $u\in\Lambda^{1}M$ implies that $\theta$ is closed. It is exact, because otherwise it is non-zero on $W$ by Poincaré duality (the de Rham cohomology of currents is equal to the de Rham cohomology of differential forms, [De]).
@@ 4.1 Barlet spaces on commendable manifolds
4.2 Non-separated points in the Teichmüller space
@@
The idea now is to take the graph $\Gamma_{t}$ of the bimeromorphism $\mathcal{X}_{t}\dashrightarrow\mathcal{M}_{t}$ and use 4.1 to show that there is a subsequence $\{\Gamma_{t_{i}}\}$ converging to a subvariety $\Gamma_{t_{0}}$ in $\mathcal{X}_{t_{0}}\times\mathcal{M}_{t_{0}}$. Then we use a topological argument to show that one of the irreducible components of $\Gamma_{t_{0}}$ is a graph of a bimeromorphism $\mathcal{X}_{t_{0}}\dashrightarrow\mathcal{M}_{t_{0}}$.
@@
Denote by $p\colon X\times M\to X$ and $q\colon X\times M\to M$ the two projections. Let us consider the action of the correspondence given by $[Z]$ on the cohomology class of the symplectic form $\Omega$. Recall that by our assumptions $[Z]_{*}[\Omega]=[\Omega^{\prime}]$. For a connected component $Z_{i}$ of the cycle $Z$ denote by $\widetilde{Z}_{i}$ its resolution of singularities and by $\widetilde{p}_{i}\colon\widetilde{Z}_{i}\to X$ and $\widetilde{q}_{i}\colon\widetilde{Z}_{i}\to M$ the induced maps. Then $[Z_{i}]_{*}[\Omega]$ is the cohomology clases of the current $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$. For $i\neq i_{0}$ the map $\widetilde{p}_{i}$ is not dominant, and since $\widetilde{q}_{i}^{*}\Omega$ is a holomorphic two-form, the current $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$ vanishes. Indeed, a closed (2,0)-current is holomorphic, hence smooth by elliptic regularity, but the support of $\widetilde{p}_{i*}(\widetilde{q}_{i}^{*}\Omega)$ is contained in a proper subvariety. Now consider the case $i=i_{0}$. Then $\widetilde{p}_{i_{0}*}(\widetilde{q}_{i_{0}}^{*}\Omega)$ is a holomorphic two-form on $X$, therefore it must be proportional to $\Omega^{\prime}$, because $X$ is hyperkähler of maximal holonomy. It follows that $\widetilde{q}_{i_{0}}^{*}\Omega$ is generically a symplectic form, so $\widetilde{q}_{i_{0}}$ must be dominant, hence $i_{0}=i_{1}$.
@@ 4.3 Degenerate twistor deformations are Kähler
Proof:  Denote by $U\subset\mathbb{A}^{1}$ the open subset of points of the base such that for $t\in U$ the fiber $\mathcal{M}_{t}$ is Kähler. Assume that $U\neq\mathbb{A}^{1}$ and let $x\in\partial U$. Then by 4.2, the fiber $\mathcal{M}_{t}$ is of Fujiki class C. Moreover, by 3 the manifold $\mathcal{M}_{t}$ admits a Hermitian symplectic form $\omega$. Then $\omega^{1,1}$ is a Hermitian metric and $dd^{c}(\omega^{1,1})=0$, therefore $\mathcal{M}_{t}$ is Kähler by [Ch, Theorem 0.2]. [^2] We thank the participants of the mathoverflow discussion for this reference It follows that $U=\mathbb{A}^{1}$.
Acknowledgements: We are grateful to Ekaterina Amerik, Alexey Golota, Evgeny Shinder, Olivier Martin and Eyal Markman for interesting discussions of the subject.